\begin{thebibliography}{23}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi

\bibitem[Barikeri et~al.(2021)Barikeri, Lauscher, Vuli{\'c}, and
  Glava{\v{s}}]{barikeri-etal-2021-redditbias}
Soumya Barikeri, Anne Lauscher, Ivan Vuli{\'c}, and Goran Glava{\v{s}}.
\newblock {R}eddit{B}ias: A real-world resource for bias evaluation and
  debiasing of conversational language models.
\newblock In \emph{Proceedings of the 59th Annual Meeting of the Association
  for Computational Linguistics and the 11th International Joint Conference on
  Natural Language Processing (Volume 1: Long Papers)}, pp.\  1941--1955,
  Online, August 2021. Association for Computational Linguistics.
\newblock \doi{10.18653/v1/2021.acl-long.151}.
\newblock URL \url{https://aclanthology.org/2021.acl-long.151}.

\bibitem[Bolukbasi et~al.(2016)Bolukbasi, Chang, Zou, Saligrama, and
  Kalai]{bolukbasi2016man}
Tolga Bolukbasi, Kai-Wei Chang, James~Y Zou, Venkatesh Saligrama, and Adam~T
  Kalai.
\newblock Man is to computer programmer as woman is to homemaker? debiasing
  word embeddings.
\newblock \emph{Advances in neural information processing systems}, 29, 2016.

\bibitem[Caliskan et~al.(2017)Caliskan, Bryson, and
  Narayanan]{caliskan2017semantics}
Aylin Caliskan, Joanna~J Bryson, and Arvind Narayanan.
\newblock Semantics derived automatically from language corpora contain
  human-like biases.
\newblock \emph{Science}, 356\penalty0 (6334):\penalty0 183--186, 2017.

\bibitem[Curtis et~al.(2019)Curtis, Jones, Tipene-Leach, Walker, Loring, Paine,
  and Reid]{curtis2019cultural}
Elana Curtis, Rhys Jones, David Tipene-Leach, Curtis Walker, Belinda Loring,
  Sarah-Jane Paine, and Papaarangi Reid.
\newblock Why cultural safety rather than cultural competency is required to
  achieve health equity: a literature review and recommended definition.
\newblock \emph{{International Journal for Equity in Health}}, 18\penalty0
  (1):\penalty0 1--17, 2019.

\bibitem[Harlow(1993)]{harlow1993}
Ray Harlow.
\newblock Lexical expansion in {Maori}.
\newblock \emph{The Journal of the Polynesian Society}, 102\penalty0
  (1):\penalty0 99--107, 1993.

\bibitem[Holtermann et~al.(2022)Holtermann, Lauscher, and
  Ponzetto]{holtermann2022fair}
Carolin Holtermann, Anne Lauscher, and Simone~Paolo Ponzetto.
\newblock Fair and argumentative language modeling for computational
  argumentation.
\newblock In \emph{Proceedings of the 60th Annual Meeting of the Association
  for Computational Linguistics (Volume 1: Long Papers)}, pp.\  7841--7861,
  2022.

\bibitem[James et~al.(2022)James, Yogarajan, Shields, Watson, Keegan, Mahelona,
  and Jones]{james-etal-2022-language}
Jesin James, Vithya Yogarajan, Isabella Shields, Catherine Watson, Peter
  Keegan, Keoni Mahelona, and Peter-Lucas Jones.
\newblock Language models for code-switch detection of te reo {M}{\=a}ori and
  {E}nglish in a low-resource setting.
\newblock In \emph{{Findings of the Association for Computational Linguistics:
  NAACL 2022}}, pp.\  650--660, Seattle, United States, July 2022. {Association
  for Computational Linguistics}.
\newblock \doi{10.18653/v1/2022.findings-naacl.49}.
\newblock URL \url{https://aclanthology.org/2022.findings-naacl.49}.

\bibitem[Jentzsch et~al.(2019)Jentzsch, Schramowski, Rothkopf, and
  Kersting]{jentzsch2019semantics}
Sophie Jentzsch, Patrick Schramowski, Constantin Rothkopf, and Kristian
  Kersting.
\newblock Semantics derived automatically from language corpora contain
  human-like moral choices.
\newblock In \emph{Proceedings of the 2019 AAAI/ACM Conference on AI, Ethics,
  and Society}, pp.\  37--44, 2019.

\bibitem[Koene et~al.(2018)Koene, Dowthwaite, and Seth]{koene2018ieee}
Ansgar Koene, Liz Dowthwaite, and Suchana Seth.
\newblock {IEEE P7003™} standard for algorithmic bias considerations: work in
  progress paper.
\newblock In \emph{Proceedings of the international workshop on software
  fairness}, pp.\  38--41, 2018.

\bibitem[Liang et~al.(2020)Liang, Li, Zheng, Lim, Salakhutdinov, and
  Morency]{liang2020towards}
Paul~Pu Liang, Irene~Mengze Li, Emily Zheng, Yao~Chong Lim, Ruslan
  Salakhutdinov, and Louis-Philippe Morency.
\newblock Towards debiasing sentence representations.
\newblock In \emph{Proceedings of the 58th Annual Meeting of the Association
  for Computational Linguistics}, pp.\  5502--5515, 2020.

\bibitem[Mahabadi et~al.(2020)Mahabadi, Belinkov, and
  Henderson]{mahabadi2020end}
Rabeeh~Karimi Mahabadi, Yonatan Belinkov, and James Henderson.
\newblock End-to-end bias mitigation by modelling biases in corpora.
\newblock In \emph{Proceedings of the 58th Annual Meeting of the Association
  for Computational Linguistics}, pp.\  8706--8716, 2020.

\bibitem[May et~al.(2019)May, Wang, Bordia, Bowman, and
  Rudinger]{may2019measuring}
Chandler May, Alex Wang, Shikha Bordia, Samuel~R Bowman, and Rachel Rudinger.
\newblock On measuring social biases in sentence encoders.
\newblock In \emph{Proceedings of NAACL-HLT}, pp.\  622--628, 2019.

\bibitem[Meade et~al.(2022)Meade, Poole-Dayan, and Reddy]{meade2022empirical}
Nicholas Meade, Elinor Poole-Dayan, and Siva Reddy.
\newblock An empirical survey of the effectiveness of debiasing techniques for
  pre-trained language models.
\newblock In \emph{Proceedings of the 60th Annual Meeting of the Association
  for Computational Linguistics (Volume 1: Long Papers)}, pp.\  1878--1898,
  2022.

\bibitem[Ravfogel et~al.(2020)Ravfogel, Elazar, Gonen, Twiton, and
  Goldberg]{ravfogel2020null}
Shauli Ravfogel, Yanai Elazar, Hila Gonen, Michael Twiton, and Yoav Goldberg.
\newblock Null it out: Guarding protected attributes by iterative nullspace
  projection.
\newblock In \emph{Proceedings of the 58th Annual Meeting of the Association
  for Computational Linguistics}, pp.\  7237--7256, 2020.

\bibitem[Rudin(2019)]{rudin2019stop}
Cynthia Rudin.
\newblock Stop explaining black box machine learning models for high stakes
  decisions and use interpretable models instead.
\newblock \emph{Nature machine intelligence}, 1\penalty0 (5):\penalty0
  206--215, 2019.

\bibitem[Schick et~al.(2021)Schick, Udupa, and Sch{\"u}tze]{schick2021self}
Timo Schick, Sahana Udupa, and Hinrich Sch{\"u}tze.
\newblock Self-diagnosis and self-debiasing: A proposal for reducing
  corpus-based bias in nlp.
\newblock \emph{Transactions of the Association for Computational Linguistics},
  9:\penalty0 1408--1424, 2021.

\bibitem[Smith et~al.(2018)Smith, Chaudhuri, Gardner, Gu, Salem, and
  L{\'e}vesque]{smith2018regulatory}
Adam~Leon Smith, Abhik Chaudhuri, Allison Gardner, Linda Gu, Malek~Ben Salem,
  and Maroussia L{\'e}vesque.
\newblock Regulatory frameworks relating to data privacy and algorithmic
  decision making in the context of emerging standards on algorithmic bias.
\newblock In \emph{{NIPS Conference Workshop on Ethical, Social and Governance
  Issues in AI, Montreal, Canada, 7th December}}, 2018.

\bibitem[Trye et~al.(2022)Trye, Yogarajan, Konig, Keegan, Bainbridge, and
  Apperley]{david-etal-2022-language}
D.~Trye, V.~Yogarajan, J.~Konig, T.~Keegan, D.~Bainbridge, and M.~Apperley.
\newblock A hybrid architecture for labelling bilingual m\={a}ori-english
  tweets.
\newblock In \emph{{Findings of the Association for Computational Linguistics:
  AACL-IJCNLP 2022}}. {Association for Computational Linguistics}, 2022.

\bibitem[Wang et~al.(2022)Wang, Zhang, and Zhu]{wang2022brief}
Xiaomeng Wang, Yishi Zhang, and Ruilin Zhu.
\newblock A brief review on algorithmic fairness.
\newblock \emph{Management System Engineering}, 1\penalty0 (1):\penalty0 7,
  2022.

\bibitem[Webster et~al.(2022)Webster, Taylor, Thomas, and
  Weller]{webster2022social}
Craig~S Webster, Saana Taylor, Courtney Thomas, and Jennifer~M Weller.
\newblock Social bias, discrimination and inequity in healthcare: mechanisms,
  implications and recommendations.
\newblock \emph{{BJA Education}}, 22\penalty0 (4):\penalty0 131--137, 2022.

\bibitem[Wilson et~al.(2022)Wilson, Tweedie, Rumball-Smith, Ross, Kazemi,
  Galvin, Dobbie, Dare, Brown, and Blakey]{wilsond-maori2022}
Daniel Wilson, Frith Tweedie, Juliet Rumball-Smith, Kevin Ross, Alex Kazemi,
  Vince Galvin, Gillian Dobbie, Tim Dare, Pieta Brown, and Judy Blakey.
\newblock {Lessons learned from developing a COVID-19 algorithm governance
  framework in Aotearoa New Zealand}.
\newblock \emph{{Journal of the Royal Society of New Zealand}}, 0\penalty0
  (0):\penalty0 1--13, 2022.
\newblock \doi{10.1080/03036758.2022.2121290}.
\newblock URL \url{https://doi.org/10.1080/03036758.2022.2121290}.

\bibitem[Yogarajan et~al.(2022)Yogarajan, Dobbie, Leitch, Keegan, Bensemann,
  Witbrock, Asrani, and Reith]{yogarajandata}
Vithya Yogarajan, Gillian Dobbie, Sharon Leitch, Te~Taka Keegan, Joshua
  Bensemann, Michael Witbrock, Varsha Asrani, and David Reith.
\newblock {Data and Model Bias in Artificial Intelligence for Healthcare
  Applications in New Zealand}.
\newblock \emph{Frontiers in Computer Science}, 4, 2022.
\newblock \doi{10.3389/fcomp.2022.1070493}.
\newblock URL
  \url{https://www.frontiersin.org/articles/10.3389/fcomp.2022.1070493}.

\bibitem[Zmigrod et~al.(2019)Zmigrod, Mielke, Wallach, and
  Cotterell]{zmigrod-etal-2019-counterfactual}
Ran Zmigrod, Sabrina~J. Mielke, Hanna Wallach, and Ryan Cotterell.
\newblock Counterfactual data augmentation for mitigating gender stereotypes in
  languages with rich morphology.
\newblock In \emph{Proceedings of the 57th Annual Meeting of the Association
  for Computational Linguistics}, pp.\  1651--1661, Florence, Italy, July 2019.
  Association for Computational Linguistics.
\newblock \doi{10.18653/v1/P19-1161}.
\newblock URL \url{https://aclanthology.org/P19-1161}.

\end{thebibliography}

\documentclass{article} 
\usepackage{iclr2023_conference_tinypaper,times}

\usepackage{amsmath,amsfonts,bm}

\def\eqref#1{equation~\ref{#1}}

\def\1{\bm{1}}

\DeclareMathAlphabet{\mathsfit}{\encodingdefault}{\sfdefault}{m}{sl}
\SetMathAlphabet{\mathsfit}{bold}{\encodingdefault}{\sfdefault}{bx}{n}

\usepackage{hyperref}
\usepackage{url}
\usepackage[para]{footmisc}
\usepackage{enumitem}   

\title{Effectiveness of Debiasing Techniques: \\ An Indigenous Qualitative Analysis}

\author{Vithya Yogarajan \& Gillian Dobbie \\
School of Computer Science \\ University of Auckland, New Zealand \\
\texttt{\{vithya.yogarajan,g.dobbie\}@auckland.ac.nz} \\
\And
Henry Gouk \\
School of Informatics, \\
University of Edinburgh, UK\\
\texttt{henry.gouk@ed.ac.uk}
}

\newcommand{\Pakeha}{P\={a}keh\={a}}
\newcommand{\Maori}{M\={a}ori}

\iclrfinalcopy 
\begin{document}

\maketitle

\begin{abstract}
An indigenous perspective on the effectiveness of debiasing techniques for pre-trained language models (PLMs) is presented in this paper. The current techniques used to measure and debias PLMs are skewed towards the US racial biases and rely on pre-defined bias attributes (e.g. ``black'' vs ``white''). Some require large datasets and further pre-training. Such techniques are not designed to capture the underrepresented indigenous populations in other countries, such as \Maori{} in New Zealand. Local knowledge and understanding must be incorporated to ensure unbiased algorithms, especially when addressing a resource-restricted society.
\end{abstract}

\section{Introduction}
The rapid increase in the use of PLMs in high-stakes decisions, such as healthcare, criminal justice and finance, have potentially profound societal implications \citep{rudin2019stop}. Studies show such PLMs capture undesirable bias and disparities; for example, a high frequency of co-occurrences in words such as ``she'' with ``cooking'' and ``he'' with ``engineer'' in training data will be reflected in PLMs  \citep{holtermann2022fair,liang2020towards,wang2022brief,schick2021self,mahabadi2020end}. While there are examples of global initiatives and legislative improvements to address the algorithmic bias \citep{koene2018ieee,smith2018regulatory}, the focus is on resource-rich large populations such as the US and Europe. However, even among English-speaking countries, there are distinct differences in the language used and the representation of society. 

This research presents a new perspective of the recently published \citet{meade2022empirical} on the effectiveness of debiasing techniques for PLMs. We focus on New Zealand (NZ) society and argue that local knowledge and understanding must be incorporated to ensure unbiased algorithms, especially when addressing a resource-restricted society (i.e. an underrepresented society with limited data).   

Aotearoa New Zealand (NZ) is a bilingual country where English is the most widely used language, and \Maori{} is the indigenous language spoken fluently by a smaller proportion of the population. NZ's unique bilingual culture is reflected in the language where loanwords from te reo \Maori{} are interlinked \citep{harlow1993,james-etal-2022-language,david-etal-2022-language}. However, the underrepresented indigenous populations, \Maori{}, experience significant inequities and social bias \citep{curtis2019cultural,webster2022social,yogarajandata}. 
We give examples of prompts (bolded) where GPT-2 generated continuations (italicised) contain evident social bias:
\par \textbf{Two brown \Maori{} men} \textit{had been seen in a car near the scene, but were not reported for more than two hours. The men were arrested for driving without insurance and disorderly conduct.}
\par \textbf{Two white kiwi men} \textit{(the kiwis are called the white people by the locals). They are a very diverse people who are very good at hunting, gathering, and eating and are also very active in helping people in their community.}

These two examples demonstrate multiple failings of GPT-2 in the context of generating continuations of NZ English text. First, there is the social bias that \Maori{} men will break the law. Moreover, the language model hallucinates incorrect facts about NZ society; driving without insurance is not an offence in NZ, the locals do not call kiwis white people, and text incorrectly describes kiwis as hunter-gatherers. 

\section{Methodology}
We consider techniques presented in \citet{meade2022empirical} for benchmarking new research in bias. Analysing debiasing techniques for PLMs requires mitigating against and measuring bias, and the model performance must be evaluated before and after debiasing. Debiasing techniques are resource-intensive; they need large datasets and further pre-training \citep{holtermann2022fair}. Moreover, post-hoc debiasing techniques such as Self-Debias cannot be evaluated against SEAT \citep{meade2022empirical}. Hence, we consider Counterfactual Data Augmentation (CDA)~\citep{zmigrod-etal-2019-counterfactual,barikeri-etal-2021-redditbias}, SentenceDebias~\citep{liang2020towards} and Iterative Nullspace Projection (INLP)~\citep{ravfogel2020null} for mitigating bias in PLMs (see Appendix~\ref{sec:debias}). 

Techniques such as StereoSet and Crowdsourced Stereotype Pairs for measuring bias are restricted by using crowd-sourced data from America or requiring hand-crafted attributes and target pairs. For NZ, adopting measures that require a large amount of data is not an option. Hence, we only consider SEAT~\citep{may2019measuring} (see Appendix~\ref{sec:seat}) with newly curated attributes and target word sets for the NZ population.

\subsection{Indigenous Prospective}
Established pre-defined lists are essential for benchmarking bias measures and techniques. However, to measure and evaluate bias in PLMs, as seen in the example presented earlier, the awareness and inclusion of indigenous perspectives is vital. Using SEAT to measure bias, and CDA, SentenceDebias, and INLP for debiasing, require pre-defined words and/or sentences \citep{meade2022empirical}. Table~\ref{tab:attribues} provides examples of commonly used pre-defined target and attribute sets, including the angry-black-woman-stereotype, Heilman-double-bind-competent-one-sentence and sent-weat6 \citep{may2019measuring,jentzsch2019semantics,meade2022empirical}.  Table~\ref{tab:attribues} also includes simple examples of NZ-specific additional attributes and target sets.  A list of bias attribute words for category {race} that capture the NZ society is also presented in Table~\ref{tab:attribues}. This research is a qualitative analysis, providing examples of attributes and data for NZ. However, it is vital to acknowledge that to quantify bias in PLMs and the effectiveness of debiasing techniques, new expert annotated baseline datasets are needed. Future directions of this research will include the establishment of such datasets.     
\vspace{-1em}
\begin{table}[t]
    \caption{Sample of target and attribute sets to measure bias (top) and bias attributes (bottom). }
    \label{tab:attribues}
    \begin{center}
    \begin{tabular}{p{0.001\textwidth}p{0.45\textwidth}|p{0.45\textwidth}}
    \hline
& \textbf{Examples of Benchmark Data} & \textbf{Examples of NZ Specific Data} \\ \hline
 1. & White \& Black Female/Male Names & Common \Maori{} and \Pakeha{} names.\\
2. & Sentences created with white or black names: eg: [Jeremiah] is a person & Sentences which also include \Maori{} names. eg: [Nikau] is a person \\
3. & Phrases such as black female, white women &  `\Maori{} woman', `\Pakeha{} female'\\ \hline
        \multicolumn{3}{c}{\bf Bias Attributes for Debiasing (category: Race)}  \\\hline
4. & (black, caucasian, asian), (black, white, china), (Africa, America, China), (Africa, Europe, Asia) & (\Maori{}, European, Asian), (\Maori{}, kiwi, Asia), (Hori, white, China), (Hori, \Pakeha{}, Asia), (Pacifica, European, Asian) \\\hline
    \end{tabular}\vspace{-1.6em}
\end{center}
\end{table}
   
\section{Discussions}
The social inequities experienced by the underrepresented indigenous population, such as \Maori{} in NZ \citep{curtis2019cultural,webster2022social,wilsond-maori2022}, is significant compared to the non-Indigenous population. As such, developing techniques for bias measures and mitigating bias must be easily adaptable towards any society, and be driven by such societies' knowledge and understanding. Using simple yet effective examples, we have outlined a vital aspect of mitigating the algorithmic bias in this new perspectives paper. Although current techniques provide a platform to understand the  bias problem, we cannot solely rely on such methods as proof that the models are unbiased. Ongoing research in mitigating bias will need to move forward from simply relying on pre-defined attribute lists or the need for large datasets. 

\subsubsection*{Acknowledgements}
VY is supported by the University of Auckland Faculty of Science Research Fellowship program and the Royal Academy of Engineering supports HG under the Research Fellowship programme.


\subsubsection*{URM Statement}
The authors acknowledge that Author VY meets the URM criteria of the ICLR 2023 Tiny Papers Track. 

\bibliography{reference}
\bibliographystyle{iclr2023_conference_tinypaper}

\appendix
\section{Appendix}

\subsection{Additional Details}
\begin{enumerate}[label=(\roman*)]
    \item Kiwi: a colloquial term for New Zealander.
    \item For common \Maori{} names please see: \url{https://www.waikato.ac.nz/library/resources/}
    \item \Pakeha{}: A non-\Maori{} New Zealander. \Pakeha{} is most commonly used to refer to white New Zealanders, but can be used to refer to anyone non-\Maori{}.
    
  \end{enumerate}

\subsection{Sentence Encoder Association Test }\label{sec:seat}
Sentence Encoder Association Test (SEAT) ~\citep{may2019measuring}, an extension to the word embedding association test (WEAT) \citep{caliskan2017semantics}, is a sentence-level bias measurement technique. In WEAT, two sets of target words $T_1$ and $T_2$ and two sets of attribute words $A_1$ and $A_2$ are used where the attribute word sets characterise bias and target word sets characterise a particular concept. WEAT measures if a word representation of a given attribute word set is more closely associated with
the representations for words from one specific
target word set. SEAT extends from WEAT by substituting the attribute words and targets words into a synthetic sentence template 
(e.g., “that is 
[WORD]”) to create a collection of sentences which can be used to measure associations. 

\subsection{Debias Techniques}\label{sec:debias}

Iterative Nullspace Projection (INLP) \citep{ravfogel2020null} is a projection-based debiasing technique that trains a linear classifier to predict the protected property we want to remove from the PLM's representation. The representations are debiased by projecting them into the nullspace of the learnt classifier's weight matrix, effectively removing all of the information the classifier used to predict the protected attribute from the representation. This process can be applied iteratively to debias the representation in PLMs. 

 SentenceDebias\citep{liang2020towards} extends a word embedding debiasing technique proposed by \citet{bolukbasi2016man} to sentence representations. 
Sentence-Debias is also a projection-based debiasing technique. By estimating a linear subspace for a particular type of bias, sentence representations are debiased by calculating the difference between the original and resulting projections of the sentence representation.   

Counterfactual Data Augmentation (CDA)~\citep{zmigrod-etal-2019-counterfactual,barikeri-etal-2021-redditbias} is a debiasing technique where a corpus is re-balanced by swapping biased attribute words, e.g. men/women. The re-balanced corpus can be used for further training to debias a modal.
 \end{document}